Consider a 342 μF capacitor charged to 1250 V in a defibrillator.

When charged to 1250 V, how much energy is stored in the capacitor? 9.60×102 J (in J)

Grade:12

1 Answers

Avishek Bhagat
3 years ago
Dear Student.
According to the question

A capacitor is a device that stores energy in the electric field between a pair of conductors on which equal but opposite electric charges have been placed.
The energy stored in a capacitor=1/2 CV2
Given,C=342μF=342×10-6F ,V=31250V
∴E=1/2×342×10-6×(1250)2= 2.67 * 102J